News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 19 July 2013 9.00am EDT

During the last week the observing schedule was replanned to include a 20 ks observation of the source GRB 130716A that was accepted as a fast turn-around Target of Opportunity on July 16. The replanned loads were uplinked on July 18.

A real-time procedure was completed on Jul 15 to dump a section of OBC-A memory as part of an acceptance test of a planned update to the Operations Control Center ground system.

Preparations are complete for the start of the summer 2013 eclipse season. The season contains 21 eclipses and runs from Jul 21 to Sep 12.

Of note this week was the announcement on Jul 12 of the Cycle 15 Peer Review recommended programs. For details see:

The schedule of targets for the next week is shown below and includes an observation of Sgr A coordinated with the EVLA.

        PSRJ1813-1246          ACIS-S       Jul 22
        M31Southeast           ACIS-S/HETG
        Radiation Belts                     Jul 23
        M31BHXN                ACIS-I       Jul 24
        M86InnerTail           ACIS-I
        TheMouse               ACIS-I       Jul 25
        Radiation Belts                     Jul 26
        HE1104-1805            ACIS-S
        LSE153                 HRC-I
        SgrA                   ACIS-S       Jul 27
        AGDra                  HRC-S/LETG
        3C289                  ACIS-S       Jul 28


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports